OpenStudy (anonymous):

Help! Calculus find the equation of the straight line tangent to the graph of the function f(x)=2x^2-7x+5 at the point (-4,65) a) where is m?, where is b?

OpenStudy (anonymous):

find f'(x) at x=-4 this is m. eq. of line is y-65=m{x-(-4)} then simplify.

OpenStudy (anonymous):

when finding f'(x) at x=-4 do i jsut sub -4 into the original equation?

OpenStudy (anonymous):

is it

OpenStudy (anonymous):

f^1(x)=4x-7

OpenStudy (anonymous):

f'(x)=4x-7 m at x=-4 is=4*-4-7=-16-7=-23 y-65=-23(x+4)